OpenStudy (anonymous):

Kenny Walks away from a blackjack table with a total of $62 in $1 and $5 chips.if kenny has 26chips,how many are $1 chips and how many are $5 chips?

OpenStudy (mathstudent55):

Let x = number of $1 chips Let y = number of $5 chips. How do you express "total of $62 in $1 and $5 chips" as an equation using x and y?

OpenStudy (anonymous):

62x-1y=5

OpenStudy (mathstudent55):

No. Be careful. The number of $1 chips is x. x number of $1 chips are worth 1 * x = x. The number of $5 chips is y. y number of $5 chips are worth 5 * y = 5y. The sum of the values is x + 5y. We know the total value is $62, so x + 5y = 62

OpenStudy (anonymous):

Oh that will be the first equation

OpenStudy (anonymous):

so x+y1=62

OpenStudy (anonymous):

sorry 1y=62

OpenStudy (mathstudent55):

Right. The second equation is simpler. There are x number of $1 chips and y number of $5 chips. How many chips are there in total?

OpenStudy (anonymous):

26chips

OpenStudy (mathstudent55):

Yes, but using x and y, how many chips are there in total?

OpenStudy (anonymous):

62+26=88chips

OpenStudy (mathstudent55):

No. 62 is dollars. 26 is number of chips. You can't add them together.

OpenStudy (anonymous):

Oh I did mistake on that

OpenStudy (mathstudent55):

We started this problem by letting x and y represent two numbers. Look above to see what x and y represent.

OpenStudy (anonymous):

we have 6chips

OpenStudy (mathstudent55):

No. We have x $1 chips and y $5 chips. The numbers of chips are x and y.

OpenStudy (anonymous):

The total chip is 26 divided by both 1 and 5 chips

OpenStudy (mathstudent55):

That's how we started solving the problem, by letting x and y represent the numbers of chips.

OpenStudy (mathstudent55):

There are x $1 chips and y $5 chips. The total number of chips is x + y We are also told there are 26 chips, so we have: x + y = 26 That's the second equation.

OpenStudy (mathstudent55):

Now we have two equations representing this problem: x + y = 26 x + 5y = 62 Now we can use the substitution method to solve it.

OpenStudy (anonymous):

x=62-5y x+y=26 x=57 57(x)+y=26 57+y=26 x=57 y=31

OpenStudy (anonymous):

Check it

OpenStudy (mathstudent55):

62 - 5y is not 57. 62 and 5y are not like terms, so they can't be combined together. x + 5y = 62 Solve for x: x = 62 - 5x Substitute in other equation: 62 - 5y + y = 26 Solve for y

OpenStudy (anonymous):

62-26=36 y=36

OpenStudy (mathstudent55):

62 - 5y + y = 26 Add -5y and y on left side: 62 - 4y = 26 Subtract 62 from both sides: -4y = -36 Divide both sides by -4: y = 9

OpenStudy (anonymous):

How did you get 4y

OpenStudy (mathstudent55):

Now that we know y, we can find x using the equation x + y = 26 x + 9 = 26 x = 17 The answer is 17 $1 chips and 9 $5 chips. Now let's check: 17 + 9 = 26 chips (Correct) 17*$1 + 9*$5 = $17 + $45 = $62 (Correct)

OpenStudy (mathstudent55):